Explain why mass is invariant under Lorentz transformations (boosts)
Thank you
1
Expert's answer
2019-09-05T09:07:00-0400

Any scalar is invariant under Lorentz transformations,

In particularly: "p_{\mu} p^{\mu}=E^2-p^2c^2= m^2c^4"

where "p_{\mu}=(E,pc)"

So, the speed of light is constant, the mass of the object is Lorentz invariant.